# Concept Steerers: Leveraging K-Sparse Autoencoders for Controllable Generations

### v1

@@ Abstract
Despite the remarkable progress in text-to-image generative models, they are prone to adversarial attacks and inadvertently generate unsafe, unethical content. Existing approaches often rely on fine-tuning models to remove specific concepts, which is computationally expensive, lack scalability, and/or compromise generation quality. In this work, we propose a novel framework leveraging k-sparse autoencoders (k-SAEs) to enable efficient and interpretable concept manipulation in diffusion models. Specifically, we first identify interpretable monosemantic concepts in the latent space of text embeddings and leverage them to precisely steer the generation away or towards a given concept (e.g., nudity) or to introduce a new concept (e.g., photographic style). Through extensive experiments, we demonstrate that our approach is very simple, requires no retraining of the base model nor LoRA adapters, does not compromise the generation quality, and is robust to adversarial prompt manipulations. Our method yields an improvement of $\mathbf{20.01\%}$ in unsafe concept removal, is effective in style manipulation, and is $\mathbf{\sim 5}$x faster than current state-of-the-art.
@@ 1 Introduction
Figure 1: Monosemantic interpretable concepts such as nudity, photographic styles, and object attributes are identified using k-sparse autoencoders (k-SAE). We leverage them to enable precise modification of a desired concept during the generation process, without impacting the overall image structure, photo-realism, visual quality, and prompt alignment (for safe concepts). Our framework can be used to remove unsafe concepts (top row), photographic styles (middle row), and object attributes (last row).
@@
Text-to-image (T2I) generative models have revolutionized content generation by producing diverse and highly photorealistic images, enabling a wide range of applications such as digital art creation (Mazzone & Elgammal, 2019), image editing (Brooks et al., 2023), and medical imaging (Kazerouni et al., 2023). These models are usually trained on several billions of web-scraped image and text pairs presumably capturing a broad spectrum of semantic concepts. Consequently, these models are also prone to be exposed to and thus generate disturbing content containing nudity, violence, child exploitation, and self-harm – raising serious ethical concerns about their downstream applications.
@@
Several attempts have been made to enforce safe generations in the past: integrating safety filters as part of the generation pipeline (Rando et al., 2022), guiding the generation process away from a pre-defined unsafe latent space (Schramowski et al., 2023), or directly erasing inappropriate concepts by modifying model weights (Gandikota et al., 2023; Heng & Soh, 2024; Li et al., 2024). While partially successful, some of these methods involve model training which is not only computationally expensive but also alters the overall model generative capabilities. More recently, a few inference-based approaches have been proposed, which do not alter model weights (Yoon et al., 2024; Jain et al., 2024). SAFREE (Yoon et al., 2024) alters the semantics of the input prompt by filtering toxic tokens, while TraSCE (Jain et al., 2024) modifies negative prompting with gradient computation to guide the model towards safer outputs. Crucially, sometimes these models have the undesirable consequence of visual degraded output generations or being misaligned with input prompts, even when the prompts are benign. Additionally, the increased inference time (e.g., $8.84$s overhead per image as noted in TraSCE (Jain et al., 2024)) due to online filtering makes them difficult to deploy in practice.
@@
In this work, we posit that the semantic information is interwoven across different layers of a generative model in complex ways that is not fully understood. Subsequently, existing training or inference-based safe generation techniques could be altering this latent landscape in undesirable ways leading to misaligned or irrelevant outputs. To this end, we approach the generation process from the ground up and explore the following crucial question: can we systematically isolate monosemantic [^1] In contrast to the one-to-many mapping of polysemantic neurons, monosemantic neurons form a one-to-one correlation with their related input features (Yan et al., 2024). concepts of varied granularities (fine-grained and abstract) from the generative latent space and surgically manipulate only them? Having such a tool would be invaluable as it would allow the user to intentionally control just the relevant concept of interest without disrupting the overall latent landscape.
@@
To this end, we leverage k-sparse autoencoders (k-SAE) (Makhzani & Frey, 2013) to design controllable generative models. k-SAEs have shown promising progress in interpreting language models by learning a sparse dictionary of monosemantic concepts (Bricken et al., 2023; Cunningham et al., 2023). In our work, we first train a k-SAE on the embeddings extracted from a corpus of text prompts containing semantic concepts we wish to control (e.g., unsafe concepts). Once trained, each k-SAE’s hidden state corresponds to an isolated monosemantic concept. During the generation process, given a concept we wish to steer, we use k-SAE to identify its corresponding latent direction and precisely manipulate the presence of that concept in the outcome, without impacting the overall generation capability (Fig. 1). Notably, our method does not require any fine-tuning as in Zhang et al. (2024), synthetic data generation as in Esposito et al. (2023), training a separate LoRA adapter (Hu et al., 2021) for each concept as in Gandikota et al. (2025) to manipulate making it fast, efficient, and adaptable to any pre-trained text to image generative framework. We summarize our empirical findings and key contributions below:
@@
  • We identify interpretable monosemantic concepts in text-to-image generation latent landscape using a k-sparse autoencoder. Once trained, k-SAE serves as a Concept Steerer to provide precise control over specific visual concepts (e.g., nudity, violence, etc.)
@@
  • Concept Steerer achieves state-of-the-art performance on unsafe concept removal while being $\sim$5x faster than the existing best method, without compromising visual quality.
@@
  • Concept Steerer is robust to adversarial prompt manipulations, achieving a 20.01% improvement against red-teaming tools, ensuring reliable image generation even under challenging scenarios.
  • Concept Steerer works out-of-the-box to any text-to-image model, is extremely simple, requires no retraining nor LoRA adapters, and is highly efficient.
@@ 2 Related Work
Controlling diffusion models: Wu et al. (2023); Wallace et al. (2024) fine-tune diffusion models using human feedback and  Bansal et al. (2023); Singhal et al. (2025) propose inference-time diffusion steering with reward functions. However, these methods rely on strong reward functions, and are computationally intensive (Uehara et al., 2025). Some methods achieve controllability by training additional modules such as low-rank adapters (LoRAs) (Gandikota et al., 2025; Stracke et al., 2025), which requires millions of parameters per concept and significantly increases generation time (Sridhar & Vasconcelos, 2024). Several inference-time intervention works attempt fine-grained control at test time. However, estimating noise at each step for each concept during generation (Brack et al., 2022, 2023) significantly slows down generation and steering model activations based on optimal transport (Rodriguez et al., 2024) requires learning activation mapping for each style. By contrast, our approach is very simple, requires no training of the base model or LoRA adapters, no additional noise/gradient computation during the generation process. Moreover, once trained, our approach allows us to manipulate any concept we want without further tuning.
@@
Safe generation: Given the growing concerns of generative models’ capability to produce inappropriate content, several valuable research has emerged in this space. Some training-based methods (Gandikota et al., 2023; Zhang et al., 2024) directly remove inappropriate concepts from the diffusion model through additional fine-tuning, while some others like  (Gandikota et al., 2024; Gong et al., 2025) update model weights to erase concepts without retraining the model. Some recently proposed inference-based approaches (Yoon et al., 2024; Jain et al., 2024) do not require training or weight updates. While effective, these methods often result in degraded image quality and increased inference time. Unlike all prior works, our method surgically isolates interpretable concepts in the generative latent space and manipulating only these in the text encoder. Thus, our approach enjoys the benefit of precise control of inappropriate concepts, does not compromise on generation quality, and maintains prompt-image alignment.
@@
Interpreting diffusion models: Recent works have demonstrated that sparse autoencoders (SAE) could recover interpretable features in large language models  (Bricken et al., 2023; Cunningham et al., 2023), CLIP vision features (Fry, 2024; Daujotas, 2024) and diffusion features (Kim et al., 2024; Surkov et al., 2024).  Kim et al. (2024) reveal monosemantic interpretable features represented within rich visual features of the diffusion model while Surkov et al. (2024) investigate how text information is integrated via cross-attention. By contrast, we focus on the text encoder of a diffusion model, identify interpretable directions via k-SAEs, and demonstrate precise steering of a variety of concepts.
@@ 3 Approach
Figure 2: K-sparse autoencoder (k-SAE) is trained on feature representations from the text encoder of the diffusion model. Once trained, it serves as a Concept Steerer, enabling precise, surgical concept manipulation by adjusting $\lambda$.
We propose a simple yet effective technique to precisely isolate and steer semantic concepts such as nudity or photographic styles using k-sparse autoencoders (Makhzani & Frey, 2013) (k-SAE). We first present how we train such a k-SAE (Sec. 3.2), followed by our method to combine different monosemantic neurons to steer abstract concepts (Sec. 3.3). We stress that a k-SAE is trained only once and no training is required for any concept the user wishes to introduce, eliminate, or modulate.
@@ 3.1 Preliminaries on text to image models
Text-to-image diffusion models (Rombach et al., 2022; Ramesh et al., 2022; Saharia et al., 2022) primarily consist of a text encoder to extract a text prompt’s intermediate embedding and a diffusion model. During training, the diffusion model progressively denoises a noisy image (or its latent representation) conditioned on the text prompt’s intermediate embedding. Formally, given an input $y_{0}$, the forward diffusion process progressively adds noise to $y_{0}$ over $T$ timesteps. The intermediate noisy image at timestep $t$ is $y_{t}=\sqrt{(1-\beta_{t})}y_{0}+\sqrt{\beta_{t}}\epsilon$ where $\epsilon$ is the Gaussian noise and $\beta_{t}$ is a timestep-dependent hyper parameter. In the reverse process, the diffusion model $\epsilon_{\theta}$ iteratively denoises $y_{t}$ at each timestep, conditioned on the text prompt embedding $c$, to predict noise $\epsilon$. The objective function for training the model is to minimize the error between the introduced and the predicted noise, defined as: $\mathbb{E}_{y,t,\epsilon\sim\mathcal{N}(0,1)}\left[\|\epsilon-\epsilon_{\theta}(y_{t},c,t)\|_{2}^{2}\right]$
@@ 3.2 Preliminaries on k-sparse autoencoders
Sparse autoencoders (Ng et al., 2011) are neural networks designed for learning compact and meaningful feature representations in an unsupervised manner. They consist of an encoder and a decoder, optimized jointly using a reconstruction loss and a sparsity regularization term to encourage only a few neurons to be maximally activated for a given input. However, the sparsity constraint introduces significant challenges during optimization (Tibshirani, 1996; Makhzani & Frey, 2013). To mitigate these issues, k-sparse autoencoders (k-SAEs) (Makhzani & Frey, 2013) were introduced. They explicitly control the number of active neurons to $k$ during training by applying a Top-$k$ activation function at each training step. Consequently, this retains only the $k$ highest activations and zeroes out the rest.
Let $W_{\text{enc}}\in\mathbb{R}^{n\times d}$ and $W_{\text{dec}}\in\mathbb{R}^{d\times n}$ represent the weight matrices of the k-SAE’s encoder and decoder respectively (Fig. 2). The hidden layer dimension $n$ is defined as an integer multiple of the input feature dimension $d$. The ratio $n/d$, referred to as the expansion factor, controls the extent to which the hidden dimension is expanded relative to the input dimension. Following Bricken et al. (2023), $b_{\text{pre}}\in\mathbb{R}^{d}$ denotes the bias term added to input $x$ before feeding to the encoder (aka pre-encoder bias), while $b_{\text{enc}}\in\mathbb{R}^{n}$ denotes the bias term of the encoder.
@@
[…]
$\small{\hat{x}=\texttt{DEC}(x)=W_{\text{dec}}z+b_{\text{pre}}},$ (2)
The training objective of a standard k-SAE is to minimize the normalized mean squared error (MSE) between the original feature $x$ and the reconstructed feature $\hat{x}$, denoted by $L_{\text{mse}}$. However, both SAEs and k-SAEs suffer from the presence of “dead latents,” where a large proportion of latents stop activating entirely at some point in training. Presence of dead latents decreases the likelihood of the network discovering separable, interpretable features while incurring unnecessary computational cost (Bricken et al., 2023). To discourage dead latents, we incorporate an auxiliary MSE loss as suggested in Gao et al. (2024). Specifically, in every training step, we identify top $k_{\text{aux}}$ dead latents and reconstruct a latent $\hat{z}$ exclusively from them, as defined below:
$\small{\hat{z}=\text{Top-}k_{\text{aux}}(\text{ReLU}(W_{\text{enc}}(x-b_{\text{pre}})+b_{\text{enc}}))},$ (3)
Now, let $\hat{e}=W_{\text{dec}}\hat{z}$ represent the reconstruction using the top $k_{\text{aux}}$ dead latents. $L_{\text{aux}}$ is defined as a reconstruction loss between the auto encoder’s residual and the output from the dead neurons ($\hat{e}$). As discussed in Gao et al. (2024), the intuition behind $L_{\text{aux}}$ is to compute gradients that push the parameters of the dead neurons in the direction of explaining the autoencoder residual ($e$). Thus, the total training loss is:
$\small{L=L_{\text{mse}}+\alpha L_{\text{aux}}=\|x-\hat{x}\|^{2}_{2}+\alpha\|e-\hat{e}\|^{2}_{2}},$ (4)

### v3

# Concept Steerers: Leveraging K-Sparse Autoencoders for Test-Time Controllable Generations

@@ Abstract
Despite the remarkable progress in text-to-image generative models, they are prone to adversarial attacks and inadvertently generate unsafe, unethical content. Existing approaches often rely on fine-tuning models to remove specific concepts, which is computationally expensive, lacks scalability, and/or compromises generation quality. In this work, we propose a novel framework leveraging k-sparse autoencoders (k-SAEs) to enable efficient and interpretable concept manipulation in diffusion models. Specifically, we first identify interpretable monosemantic concepts in the latent space of text embeddings and leverage them to precisely steer the generation away or towards a given concept (e.g., nudity) or to introduce a new concept (e.g., photographic style) – all during test time. Through extensive experiments, we demonstrate that our approach is very simple, requires no retraining of the base model nor LoRA adapters, does not compromise the generation quality, and is robust to adversarial prompt manipulations. Our method yields an improvement of $\mathbf{20.01\%}$ in unsafe concept removal, is effective in style manipulation, and is $\mathbf{\sim 5}$x faster than the current state-of-the-art. Code is available at: https://github.com/kim-dahye/steerers
@@ 1 Introduction
Text-to-image (T2I) generative models have revolutionized content generation by producing diverse and highly photorealistic images, enabling a wide range of applications such as digital art creation mazzone2019art , image editing brooks2023instructpix2pix , and medical imaging kazerouni2023diffusion . These models are usually trained on several billions of web-scraped image and text pairs presumably capturing a broad spectrum of semantic concepts. Consequently, these models are also prone to be exposed to and thus generate disturbing content containing nudity, violence, child exploitation, and self-harm – raising serious ethical concerns about their downstream applications.
Figure 1: Monosemantic interpretable concepts such as nudity, photographic styles, and object attributes are identified using k-sparse autoencoders (k-SAE). We leverage them to enable precise modification of a desired concept during the generation process, without impacting the overall image structure, photo-realism, visual quality, and prompt alignment (for safe concepts). Our framework can be used to remove unsafe concepts (left), photographic styles (middle), and object attributes (right).
@@
Several attempts have been made to enforce safe generations in the past: integrating safety filters as part of the generation pipeline rando2022red , guiding the generation process away from a pre-defined unsafe latent space schramowski2023safe , or directly erasing inappropriate concepts by modifying model weights gandikota2023erasing ; heng2024selective ; li2024safegen . While partially successful, some of these methods involve model training which is not only computationally expensive but also alters the overall model’s generative capabilities. More recently, a few inference-based approaches have been proposed, which do not alter model weights yoon2024safree ; jain2024trasce . SAFREE yoon2024safree alters the semantics of the input prompt by filtering toxic tokens, while TraSCE jain2024trasce modifies negative prompting with gradient computation to guide the model towards safer outputs. Crucially, sometimes these models have the undesirable consequence of visually degraded output generations or being misaligned with input prompts, even when the prompts are benign. Additionally, the increased inference time (e.g., $8.84$s overhead per image as noted in TraSCE jain2024trasce ) makes them difficult to deploy in practice.
@@
In this work, we posit that the semantic information is interwoven across different layers of a generative model in complex ways that is not fully understood. Subsequently, existing training or inference-based safe generation techniques could be altering this latent landscape in undesirable ways leading to misaligned or irrelevant outputs. To this end, we approach the generation process from the ground up and explore the following crucial question: can we systematically isolate monosemantic concepts of varied granularities (fine-grained and abstract) from the generative latent space and surgically manipulate only them? [^1] In contrast to the one-to-many mapping of polysemantic neurons, monosemantic neurons form a one-to-one correlation with their related input features yan2024encourage . Such an invaluable tool would allow the user to intentionally control just the relevant concept of interest without disrupting the overall latent landscape.
@@
To this end, we leverage k-sparse autoencoders (k-SAE) makhzani2013k to design controllable generative models. k-SAEs have shown promising progress in interpreting language models by learning a sparse dictionary of monosemantic concepts bricken2023monosemanticity ; sparse_lang . In our work, we first train a k-SAE on the embeddings extracted from a corpus of text prompts containing semantic concepts we wish to control (e.g., unsafe concepts). Once trained, each k-SAE’s hidden state corresponds to an isolated monosemantic concept. During the generation process, given a concept we wish to steer, we use k-SAE to identify its corresponding latent direction and precisely manipulate the presence of that concept in the outcome, without impacting the overall generation capability (Fig. 1). Notably, our method does not require any fine-tuning as in zhang2024forget , synthetic data generation as in esposito2023mitigating , training a separate LoRA adapter hu2021lora for each concept as in gandikota2025concept to manipulate making it fast, efficient, and adaptable to any pre-trained text to image generative framework. We summarize our findings and key contributions below:
@@
  • We identify interpretable monosemantic concepts in text-to-image generation latent landscape using a k-sparse autoencoder. The k-SAE serves as a Concept Steerer and offers precise control over semantic concepts (e.g., nudity, photographic style, etc.)
@@
  • Concept Steerer achieves state-of-the-art performance on unsafe concept removal while being $\sim$5x faster than the existing best method jain2024trasce , without compromising visual quality.
@@
  • Concept Steerer is robust to adversarial prompt manipulations, achieves a 20.01% improvement against red-teaming algorithms, ensuring reliable image generation even under challenging scenarios.
  • Concept Steerer works out-of-the-box at test-time, can be applied to any text-to-image generative model, requires no retraining nor LoRA adapters, is simple and efficient.
@@ 2 Related Work
Controlling diffusion models: wu2023human ; wallace2024diffusion fine-tune diffusion models using human feedback and  bansal2023universal ; singhal2025general propose inference-time diffusion steering with reward functions. However, these methods rely on strong reward functions, and are computationally intensive uehara2025reward . Some methods achieve controllability by training additional modules such as low-rank adapters (LoRAs) gandikota2025concept ; stracke2025ctrloralter , which requires millions of parameters per concept and significantly increases generation time sridhar2024prompt . Several inference-time intervention works attempt fine-grained control at test time. However, estimating noise at each step for each concept during generation brack2022stable ; brack2023sega significantly slows down generation and steering model activations based on optimal transport rodriguez2024controlling requires learning activation mapping for each style. By contrast, our approach is very simple, requires no training of the base model or LoRA adapters, no additional noise/gradient computation during the generation process. Moreover, once trained, our approach allows us to manipulate any concept we want without further tuning. To the best of our knowledge, ours is the only work which leverages sparse autoencoders to offer more creative and generative control to users.
@@
Safe generation: Given the growing concerns of generative models’ capability to produce inappropriate content, several valuable research has emerged in this space. Some training-based methods gandikota2023erasing ; heng2024selective ; zhang2024forget ; kumari2023ablating ; lu2024mace ; fan2023salun ; bui2024erasing ; ko2024boosting ; hong2024all ; park2024direct ; lyu2024one ; huang2024receler ; kim2024safeguard ; chen2024score ; chavhan2024conceptprune directly remove inappropriate concepts from the diffusion model through additional fine-tuning, while some others like  gandikota2024unified ; gong2025reliable update model weights to erase concepts without retraining the model. Some recently proposed inference-based approaches yoon2024safree ; jain2024trasce do not require training or weight updates. While effective, these methods often result in degraded image quality and increased inference time. Unlike all prior works, our method surgically isolates interpretable concepts in the generative latent space and manipulating only these in the text encoder. Thus, our approach enjoys the benefit of precise control of inappropriate concepts, does not compromise on generation quality, and maintains prompt-image alignment. In the similar vein, SDID li2024self discovers interpretable latent directions for a given concept in the bottleneck of the diffusion model. However, it requires generating $1000$ paired unsafe and safe images to learn a representation for each unsafe concept (e.g., 1000 pairs of images of violence and non-violence), making this approach less flexible and scalable. By contrast, our approach does not require thousands of generating pairs of images per concept. Instead, once Concept Steerer is trained on a set of prompts, it can steer various concepts and can be applied to different text-to-image models.
@@
Interpreting diffusion models: Recent works have demonstrated that sparse autoencoders (SAE) could recover interpretable features in large language models  bricken2023monosemanticity ; sparse_lang , CLIP vision features fry2024towards ; daujotas2024interpreting , multimodal LLMs pach2025sparse ; yan2025multi and diffusion features kim2024textit ; surkov2024unpacking .  kim2024textit reveals monosemantic interpretable features represented within rich visual features of the diffusion model while surkov2024unpacking investigates how text information is integrated via cross-attention. However, in all prior works, the interpretation is done through manual inspection and/or visualization, which limits their scalability. By contrast, Concept Steerer offers an automated way to interpret the concepts it is trained on through steering the joint latent space of diffusion models.
@@ 3 Approach
Figure 2: K-sparse autoencoder (k-SAE) is trained on feature representations from the text encoder of the diffusion model. Once trained, it serves as a Concept Steerer, enabling precise concept manipulation at test-time. $\lambda$ denotes the strength of the concept.
We propose a simple yet effective technique to precisely isolate and steer semantic concepts such as nudity or photographic styles using k-sparse autoencoders makhzani2013k (k-SAE). We first present how we train such a k-SAE (Sec. 3.2) on text representations, followed by our method to combine different monosemantic neurons to steer abstract concepts (Sec. 3.3). We stress that a k-SAE is trained only once and no training is required for any concept the user wishes to introduce, eliminate, or modulate.
@@ 3.1 Preliminaries on text to image models
Text-to-image diffusion models rombach2022high ; ramesh2022hierarchical ; saharia2022photorealistic primarily consist of a text encoder to extract a text prompt’s intermediate embedding and a diffusion model. During training, the diffusion model progressively denoises a noisy image (or its latent representation) conditioned on the text prompt’s intermediate embedding. Formally, given an input $y_{0}$, the forward diffusion process progressively adds noise to $y_{0}$ over $T$ timesteps. The intermediate noisy image at timestep $t$ is $y_{t}=\sqrt{(1-\beta_{t})}y_{0}+\sqrt{\beta_{t}}\epsilon$ where $\epsilon$ is the Gaussian noise and $\beta_{t}$ is a timestep-dependent hyper parameter. In the reverse process, the diffusion model $\epsilon_{\theta}$ iteratively denoises $y_{t}$ at each timestep, conditioned on the text prompt embedding $c$, to predict noise $\epsilon$. The objective function for training the model is to minimize the error between the introduced and the predicted noise, defined as: $\mathbb{E}_{y,t,\epsilon\sim\mathcal{N}(0,1)}\left[\|\epsilon-\epsilon_{\theta}(y_{t},c,t)\|_{2}^{2}\right].$
@@ 3.2 Preliminaries on k-sparse autoencoders
Sparse autoencoders ng2011sparse are neural networks designed for learning compact and meaningful feature representations in an unsupervised manner. They consist of an encoder and a decoder, optimized jointly using a reconstruction loss and a sparsity regularization term to encourage only a few neurons to be maximally activated for a given input. However, the sparsity constraint introduces significant challenges during optimization makhzani2013k ; tibshirani1996regression . To mitigate these issues, k-sparse autoencoders (k-SAEs) makhzani2013k were introduced. They explicitly control the number of active neurons to $k$ during training by applying a Top-$k$ activation function at each training step. Consequently, this retains only the $k$ highest activations and zeroes out the rest.
Let $W_{\text{enc}}\in\mathbb{R}^{n\times d}$ and $W_{\text{dec}}\in\mathbb{R}^{d\times n}$ represent the weight matrices of the k-SAE’s encoder and decoder respectively (Fig. 2). The hidden layer dimension $n$ is defined as an integer multiple of the input feature dimension $d$. The ratio $n/d$, referred to as the expansion factor, controls the extent to which the hidden dimension is expanded relative to the input dimension. Following bricken2023monosemanticity , $b_{\text{pre}}\in\mathbb{R}^{d}$ denotes the bias term added to input $x$ before feeding to the encoder (aka pre-encoder bias), while $b_{\text{enc}}\in\mathbb{R}^{n}$ denotes the bias term of the encoder.
@@
where the $\text{Top-}k$ function retains only the top $k$ neuron activations and sets the remaining activations to zero makhzani2013k . The decoder reconstructs $\hat{x}$ as:
$\small{\hat{x}=\texttt{DEC}(x)=W_{\text{dec}}z+b_{\text{pre}}}.$ (2)
The training objective of a standard k-SAE is to minimize the normalized mean squared error (MSE) between the original feature $x$ and the reconstructed feature $\hat{x}$, denoted by $L_{\text{mse}}$. ReLU and Top-K activation functions in Eq. 1 ensure that k-SAEs do not learn a linear mapping between inputs and outputs. However, both SAEs and k-SAEs suffer from the presence of “dead latents,” where a large proportion of latents stop activating entirely at some point in training. Presence of dead latents decreases the likelihood of the network discovering separable, interpretable features while incurring unnecessary computational cost bricken2023monosemanticity . To discourage dead latents, we incorporate an auxiliary MSE loss as suggested in gao2024scaling . Specifically, in every training step, we identify top $k_{\text{aux}}$ dead latents (i.e., latents that have the least activation value) and reconstruct a latent $\hat{z}$ exclusively from them, as defined below:
$\small{\hat{z}=\text{Top-}k_{\text{aux}}(\text{ReLU}(W_{\text{enc}}(x-b_{\text{pre}})+b_{\text{enc}}))}.$ (3)
Now, let $\hat{e}=W_{\text{dec}}\hat{z}$ represent the reconstruction using the top $k_{\text{aux}}$ dead latents. $L_{\text{aux}}$ is defined as a reconstruction loss between the auto encoder’s residual and the output from the dead neurons ($\hat{e}$). As discussed in gao2024scaling , the intuition behind $L_{\text{aux}}$ is to compute gradients that push the parameters of the dead neurons in the direction of explaining the autoencoder residual ($e$). Thus, the total training loss is:
$\small{L=L_{\text{mse}}+\alpha L_{\text{aux}}=\|x-\hat{x}\|^{2}_{2}+\alpha\|e-\hat{e}\|^{2}_{2}}.$ (4)